OpenStudy (anonymous):

What is the equation in standard form of a perpindicular line that passes through (5,1)

OpenStudy (anonymous):

perpendicular to what?

OpenStudy (anonymous):

I don't know it passes through 5.1 though. Algebra is confusing

OpenStudy (anonymous):

there are an infinite number of lines passing through \((5,1)\) i am sure there is another part to this question perpendicular to the line blah blah perpendicular to the x axis perpendicular to the y axis perpendicular to something "perpendicular" means two lines make a right angle, so it is a comparison to something

OpenStudy (anonymous):

Is there a slope given?

OpenStudy (anonymous):

Nope that was the full question. So far I got y-1=m(x-5)

OpenStudy (anonymous):

Are you sure there isn't any other information given? An equation? A line?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

y = (-1/m)(x-5) + 1, given the equation of the original line is y = mx + b. Simply just take the negative reciprocal of m for the slope.

OpenStudy (anonymous):

Ok ya so confused.

OpenStudy (anonymous):

Then, (y-1)/(x-5) = -1/m. Rearrange into slope-intercept form. Now, you're done!

OpenStudy (anonymous):

If you're still confused, just ask a question and maybe I can guide you through it.

OpenStudy (anonymous):

Well the next ones ants the x intercept of it and that's gonna be kindofhard to figure for me.

OpenStudy (anonymous):

Just take the answer, y = (-1/m)(x-5) + 1, set y=0, and solve for x. The x-intercept takes the form (a,0). You've got the equation now. The value for x will be expressed in terms of m.

OpenStudy (anonymous):

ax + by +c and wq of a line perpendicular to it is bx - ay + c

OpenStudy (anonymous):

Note : Product of their Slopes m1*m2 = -1

OpenStudy (anonymous):

You see, your original question never gave the equation of the "original" line. So, if the question is complete and is to make sense, it logically follows that they are talking about a line in general. So, you have the leeway to say that whatever line you are taling about in the beginning, it has the form of y=mx + b. And that's how I derived the answer.

OpenStudy (anonymous):

If you are still confused, just ask a question and I'll walk you through it.